# Periodic orbits and their gravitational waves in EMRIs: supermassive black hole affected by galactic dark matter halos

### v1

@@ Appendix A Periodic Orbits under the Condition of Selected Energy $E$
The periodic orbits are determined by the orbit parameters $(E,L)$. In Section III, we characterize periodic orbits by fixing the angular momentum at $L=L_{\rm ISCO}+\varepsilon(L_{\rm MBO}-L_{\rm ISCO})$ and varying the energy $E_{(z~w~v)}$ across different zoom-whirl-vertex integers (zwv)(z~w~v). In this Appendix, we adopt an alternative but equivalent approach: we fix the energy parameter $E$ and vary the angular momentum $L_{(z~w~v)}$ to explore different periodic orbit patterns (zwv)(z~w~v).
@@
To perform our analysis, we set the energy to $E=0.96$ to study how dark matter mass affects the precession angle. Other choices of energy values lead to the same conclusions. The relationship between the precession angle $q$ and the angular momentum $L$ for different dark matter masses $k$ is shown in Figure 20. The numerical results indicate that for any given precession angle $q$, increasing the dark matter mass $k$ requires higher angular momentum $L$. This is because the dark matter mass alters the gravitational potential, which increases the angular momentum needed to achieve a particular orbital precession. Fig. 21 provides detailed precession characteristics for each model at fixed dark matter masses. An important observation is that the precession angle curves for the NFW and Beta models overlap completely, even at large dark matter masses. For the same precession angle $q$, both NFW and Beta models require higher angular momentum $L$ than the Moore model.
Based on this precession angle analysis, we present the corresponding periodic orbits shown in Fig. 22; their angular momentum values are listed in Table 3. The overall patterns in orbital shape evolution follow those observed in the fixed angular momentum case (in Fig. 6). The orbital paths predicted by the NFW and Beta models remain identical for all dark matter masses $k$, which aligns with their coinciding precession angle curves in Fig. 21. The influence of dark matter becomes increasingly evident as $k$ grows: orbits in the presence of dark matter halos diverge more substantially from those around a Schwarzschild black hole. We observe that higher dark matter masses systematically increase the apoapsis distance, making the orbital shapes gradually more distinct from Schwarzschild orbits. These results are consistent with the conclusions in Section III.1, where we analyzed periodic orbits using angular momentum $L=L_{\text{ISCO}}+\varepsilon(L_{\text{MBO}}-L_{\text{ISCO}})$.
@@ Appendix B The Parameter $E_{(z~w~v)}$ and $L_{(z~w~v)}$ for Periodic Orbits Around Black Holes
As established in Section III, the characteristics of a periodic orbit, identified by its zoom-whirl-vertex numbers (zwv)(z~w~v), are determined by its energy $E_{(z~w~v)}$ and angular momentum $L_{(z~w~v)}$. In Sections III.1 and III.2, we systematically analyze the influence of the dark matter halo by varying the mass parameter $k$ (from $1\times 10^{3}M\sim 2\times 10^{4}M$) and the scale parameter $h$ (from $10^{7}M\sim 10^{10}M$), while keeping the angular momentum fixed at $L$ (with $\varepsilon=0.5$ set in Eq. (24). The corresponding energy values $E_{(z~w~v)}$ for the periodic orbits shown in Figs. 6 and 10 are listed in Tables 1 and 2, respectively. Furthermore, an alternative analysis is conducted in Appendix B with a fixed energy of $E=0.96$. For this case, Table 3 provides the angular momentum values $L_{(z~w~v)}$ for the periodic orbits plotted in Fig. 22 for various dark matter models and different values of the dark matter mass parameter $k$.

### v2

@@ Appendix A Periodic Orbits under the Condition of Selected Energy $E$
The periodic orbits are determined by the orbit parameters $(E,L)$. In Section III, we characterize periodic orbits by fixing the angular momentum at $L=L_{\rm ISCO}+\varepsilon(L_{\rm MBO}-L_{\rm ISCO})$ and varying the energy $E_{(z~w~v)}$ across different zoom-whirl-vertex integers (zwv)(z~w~v). In this Appendix, we adopt an alternative but equivalent approach: fixing the energy parameter $E$ and obtaining the angular momentum $L_{(z~w~v)}$ to explore different periodic orbit configurations (zwv)(z~w~v).
@@
To perform our analysis, we set the energy to $E=0.96$ to study how dark matter mass affects the precession angle. Other choices of energy values lead to the same conclusions, which are not presented in the present work. The relationship between the precession angle $q$ and the angular momentum $L$ for different dark matter masses $k$ is shown in Figure 20. The numerical results indicate that for any given precession angle $q$, increasing the dark matter mass $k$ requires higher angular momentum $L$. This is because the dark matter mass alters the effective potential of test particles, which increases the angular momentum needed to achieve a particular orbital precession. Fig. 21 provides detailed precession characteristics for each model at fixed dark matter masses. An important observation is that the precession angle curves for the NFW and Beta models overlap completely, even at large dark matter masses. For the same precession angle $q$, both NFW and Beta models require higher angular momentum $L$ than the Moore model.
Based on this precession angle analysis, we present the corresponding periodic orbits (shown in Fig. 22); their angular momentum values are listed in Table 3. The overall patterns and characteristics in orbital shape with different orbital configurations (zwv)(z~w~v) follow those observed in the fixed angular momentum case (Fig. 6). Besides, the periodic orbits calculated in the NFW and Beta models are relatively more extended than those in the Moore model. The orbital trajectories predicted by the NFW and Beta models remain identical for all dark matter masses $k$, which aligns with their coinciding precession angle curves in Fig. 21. The influence of dark matter becomes increasingly evident as $k$ grows: periodic orbits around SMBH in the presence of dark matter halos diverge more substantially from those around a Schwarzschild black hole with a increasement in the apoapsis distance. These results are consistent with the conclusions in Section III.1, where we analyzed periodic orbits using the angular momentum $L=L_{\text{ISCO}}+\varepsilon(L_{\text{MBO}}-L_{\text{ISCO}})$.
@@ Appendix B The Parameter $E_{(z~w~v)}$ and $L_{(z~w~v)}$ for Periodic Orbits Around Black Holes
As established in Section III, the characteristics of a periodic orbit, identified by its zoom-whirl-vertex numbers (zwv)(z~w~v), are determined by its energy $E_{(z~w~v)}$ and angular momentum $L_{(z~w~v)}$. In Sections III.1 and III.2, we systematically analyze the influence of the dark matter halo by varying the mass parameter $k$ (from $1\times 10^{3}M\sim 2\times 10^{4}M$) and the scale parameter $h$ (from $10^{7}M\sim 10^{10}M$), while keeping the angular momentum fixed at $L=L_{\text{ISCO}}+\varepsilon(L_{\text{MBO}}-L_{\text{ISCO}})$ with $\varepsilon=0.5$ selected as a representative case to present numerical results on periodic orbits. The corresponding energy values $E_{(z~w~v)}$ for the periodic orbits shown in Figs. 6 and 10 are listed in Tables 1 and 2, respectively. Furthermore, an alternative analysis is conducted in Appendix B with a selected energy of $E=0.96$. For this case, Table 3 provides the angular momentum values $L_{(z~w~v)}$ for the periodic orbits plotted in Fig. 22 for various dark matter models and different dark matter mass parameter $k$.